OpenStudy (jewelie):

2x+8y=12 what is the slope of the graph of this equation?

OpenStudy (festinger):

Can you make y the subject of the equation?

OpenStudy (jewelie):

What do you mean?

OpenStudy (lowkey.s):

WELCOME TO OPENSTUDY!

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)

likeabossssssss (likeabossssssss):

what do u think it is ?

OpenStudy (lowkey.s):

2x+8y=12 8y=12-2x y=3/2 - 1x/4

OpenStudy (festinger):

Can you convert the equation from\[2x+8y=12\] into something of the form \[y= mx+c\]

OpenStudy (jewelie):

It can be converted into ax+yb=c

OpenStudy (lowkey.s):

Like what @festinger

OpenStudy (lowkey.s):

like what he says

OpenStudy (lowkey.s):

*

OpenStudy (jewelie):

Do one of you guys have a social media I can show you a picture of my work so we both understand?

OpenStudy (holly345):

In the equation y=mx+b, where m is the slope of the equation, m=-1.